5x^2-11x+2=0
my answer is (5x-1)(x-2)

Answer by dabanfield(803)

5x^2-11x+2=0
my answer is (5x-1)(x-2)
This is correct so far. To find the solutions for the original equation you need to set each factor equal to zero(remember that a*b = 0 if and only if a = 0 and/or b = 0).
So you need to solve for x in:
5x - 1 = 0
and
x - 2 = 0
These two values of x are the solutions for the original equation.